OpenStudy (anonymous):

PLEASE HELP ILL GIVE OUT A MEDAL AND BEST ANSWER what is the molar mass of A1(C1O2)3 ? Explain how you calculated this value.

thomaster (thomaster):

Well that compound does not exist... however I think you confused the 1 with the lowercase letter L. So it would be \(\Large Al(ClO_2)_3\) which is the formula for aluminium chlorite. To calculate the molar mass, you check how many atoms of the same element you got. We got 1 aluminum atom the part in parenthesis is multiplied by 3 because of the subscript 3 after the parenthesis. So we have 3 chlorine atoms and 2*3= 6 oxygen atoms Now we go to a periodic table ( http://ptable.com ) and check for the atomic mass of each element. Aluminium = 26.981 Chlorine = 35.45 Oxygen = 15.999 Now just multiply that with the amount of atoms you have. So: Al = 1 * 26.981 = 26.981 Cl = 3 * 35.45 = 106.35 O = 6 * 15.999 = 95.994 we add that ------ + And we have: 229.325 g/mol

OpenStudy (anonymous):

THANK YOU SO MUCHHH :)